# HtmlRAG: HTML is Better Than Plain Text for Modeling Retrieved Knowledge in RAG Systems

### v1

@@ Abstract.
Retrieval-Augmented Generation (RAG) has been shown to improve knowledge capabilities and alleviate the hallucination problem of LLMs. The Web is a major source of external knowledge used in RAG systems, and many commercial systems such as ChatGPT and Perplexity have used Web search engines as their major retrieval systems. Typically, such RAG systems retrieve search results, download HTML sources of the results, and then extract plain texts from the HTML sources. Plain text documents or chunks are fed into the LLMs to augment the generation. However, much of the structural and semantic information inherent in HTML, such as headings and table structures, is lost during this plain-text-based RAG process. To alleviate this problem, we propose HtmlRAG, which uses HTML instead of plain text as the format of retrieved knowledge in RAG. We believe HTML is better than plain text in modeling knowledge in external documents, and most LLMs possess robust capacities to understand HTML. However, utilizing HTML presents new challenges. HTML contains additional content such as tags, JavaScript, and CSS specifications, which bring extra input tokens and noise to the RAG system. To address this issue, we propose HTML cleaning, compression, and pruning strategies, to shorten the HTML while minimizing the loss of information. Specifically, we design a two-step block-tree-based pruning method that prunes useless HTML blocks and keeps only the relevant part of the HTML. Experiments on six QA datasets confirm the superiority of using HTML in RAG systems  [^1] Code and datasets are available at https://github.com/plageon/HtmlRAG..
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: TheWebConf 2025; June 03–05, 2025; Woodstock, NYisbn: 978-1-4503-XXXX-X/25/06ccs: Information systems Web search engines
@@ 1. Introduction
However, employing HTML as the knowledge format for LLMs also presents the challenge of handling longer input sequences and noisy contexts. Our preliminary experiments show that a real HTML document from the Web contains over 80K tokens on average, among which over 90% of the tokens are CSS styles, JavaScript, Comments, or other meaningless tokens. Compared to the common maximum context window of current LLMs, which ranges from 32K to 128K, an individual document length of 80K is unacceptable. The noisy tokens. The aforementioned meaningless tokens in HTML documents can also affect the generation quality of LLMs. To solve this problem, in this paper, we devise a HTML Cleaning module to remove semantically irrelevant content in HTML documents, while keeping the main content intact. We also adjust the HTML tree structure without losing semantic information, for example, merging multiple layers of single nested HTML tags and removing empty tags. These processes reduce the length of the HTML to 6% of its original size.
Even after cleaning, HTML documents remain relatively long (over 4K each) to LLMs. To shorten the input context and remove the noise contained in the original retrieved documents, existing RAG systems have utilized different types of post-retrieval result refiners (Zhou et al., 2024b; Jin et al., 2024b; Xu et al., 2024; Jiang et al., 2024). These refiners extract the relevant text chunks or key sentences from the documents, regarding the user’s query and LLMs’ preference, and discard other content. These plain-text-based refiners cannot be directly applied to HTML because simply chunking HTML without considering its structure may generate unreasonable chunks. Hence, we further design an HTML Pruning module, which functions upon the intrinsic tree structure of HTML. The pruning process is comprised of the following steps:
@@ 2.1. Retrieval-Augmented Generation (RAG)
RAG systems augment LLM with external knowledge. A typical RAG pipeline includes components such as a query rewriter (Tan et al., 2024), a retriever (Li et al., 2024; Shi et al., 2024), a reranker (Wang et al., 2024c; Shi et al., 2024), a refiner (Xu et al., 2024; Jin et al., 2024b; Jiang et al., 2024), and a reader (Zhu et al., 2024; Asai et al., 2024). This typical pipeline is widely used by mainstream RAG frameworks, such as LangChain (Chase, 2022) and LlamaIndex (Liu, 2022). Many works aim to optimize components in the pipeline, and previous works also manage to enhance the performance of RAG in other ways. Some methods devise new RAG frameworks, like retrieving external knowledge actively when internal knowledge is missing (Tan et al., 2024; Jiang et al., 2023b; Asai et al., 2024), or letting the LLM plan the retrieval process in a straight line or a tree structure (Shao et al., 2023; Kim et al., 2023b). However, most existing RAG systems take plain text as the format of external knowledge. Instead, we propose to take HTML as the format of external knowledge, and we believe using HTML can keep richer semantics in retrieved results.
@@ 2.2. Post-Retrieval Process of RAG
RAG systems usually apply post-retrieval processes (i.e., result refiners) to extract only the useful content to shorten the input context sent to LLMs. The chunking-based refiner is a widely used solution, which first chunks the text according to certain rules, and then uses a reranking model to select top chunks with high relevance (Karpukhin et al., 2020; Moniz et al., 2024). Another solution is abstractive refiner, which utilizes a text-to-text language model to generate abstracts of results (Xu et al., 2024; Jiang et al., 2024; Gilbert et al., 2023). Some works use off-the-shelf abstractive models (Zhang et al., 2023a, b) or fine-tuned abstractive models (Jiang et al., 2024) to summarize retrieved results in a segmented and hierarchical manner. Others leverage the logits of language models to determine the importance of words within documents (Liu, 2019; Li, 2023).
@@
Figure 2. HTML for RAG pipeline overview
@@
Overview of the pipeline in HtmlRAG
@@ 3.2. HTML Cleaning
Since the original HTML documents are excessively long (over 80K each), and it’s needless to involve semantic features, model-based methods are inappropriate at this step. Thus, we first design a rule-based HTML cleaning, which pre-processes the HTML without considering the user’s query. This cleaning process removes irrelevant content and compresses redundant structures, retaining all semantic information in the original HTML. The compressed HTML of HTML cleaning is suitable for RAG systems equipped with long-context LLMs and are not willing to lose any information before generation. The cleaned HTML also serves as the basis for the following HTML pruning.
@@ 3.2.2. Lossless Structural Compression
3.3. Granularity-Adjustable Block Tree Construction
@@
Considering the above problem, we propose an optimized tree structure that models HTML, which is not so fine-grained. Ideally, the granularity of the tree structure can be adjusted for different pruning requirements. We term it as a “block tree”, and we set the maximum number of words per block, $maxWords$ to control the granularity of the block tree. In terms of block tree construction, we start from a DOM tree, and we merge fragmented child nodes into their parent and treat them as a block. We can recursively merge blocks or child nodes into their parent to form a bigger block under the condition that the number of words in a block does not exceed $maxWords$. After merging, original leaf nodes that are unable to be merged are also regarded as blocks. Algorithm details are demonstrated in Appendix B.
@@ 3.4.1. Pruning Blocks based on Text Embedding
The refining process is expected to shorten the retrieval results while preserving key information as much as possible. A straightforward idea is to extract plain text in the block and calculate a similarity score with the user’s query using text embeddings. Then we use a greedy algorithm to prune the block tree by deleting low-similarity blocks and retraining higher ones. In practice, we keep deleting the block with the lowest relevance until the total length of the HTML documents satisfies the context window we set. After block deleting, redundant HTML structures will re-appear, so we re-adjust the HTML structure, meaning multiple layers of single-nested tags are merged and empty tags are removed. The detailed pruning algorithm is demonstrated in Appendix B.
@@
3.4.2. Generative Fine-Grained Block Pruning
To further prune blocks with a finer granularity, we expand the leaf nodes of the pruned block tree and get a finer-grained block tree. Given the limitations of the embedding-model-based block pruning, we propose to use a generative model because it has a long context to cover the whole block tree and is not limited to modeling one block at a time. Yet processing the cleaned HTML directly with a generative model is inappropriate because the cleaned HTML is long (60K on average), which brings much computational cost. Similarly, the generative model is supposed to calculate scores for blocks. Inspired by CFIC (Qian et al., 2024), which takes the text chunk’s sequence generation probability as the score for that chunk, we propose to use a sequence of tags to identify a block. Specifically, the sequence consists of tags starting from the root tag and walking down to the block’s tag, and we term this sequence as “block path”. In the inference phase, the generative model follows the structure of the block tree and calculates the scores of blocks in the block tree. The scores of blocks are derived from the token logits, as displayed in Figure 3. At last, we use the same block pruning operation as we mention in §3.4.1 to obtain the refined HTML document.
@@ 3.4.2. Generative Fine-Grained Block Pruning
We collect a small amount of supervised data to enhance the model’s capability in block path generation. Following the typical SFT process (Qin et al., 2024), the steps for training data collecting, filtering, and constructing are as follows: First, we sample queries from the training set of several open-source QA datasets. For each query, we retrieve a couple of related HTML documents using the online search engine Bing. Then we clean the retrieved HTML, and prune the HTML with the embedding model. By adjusting the output length in HTML pruning, we get pruned HTML documents of various lengths, ranging from 2K tokens to 32K tokens. After that, we build a block tree from each HTML document pruned by the embedding model, and calculate the exact match score for the content within blocks with the gold answer. To ensure the data quality, we discard samples in which no block’s content exactly matches the gold answer, meaning highly relevant HTML documents are not retrieved. More training details are discussed in Appendix A.
@@
In the first two conditions, it is needless to infer with the generative model, meaning many tokens can be skipped. This brings down the inference computational cost. Apart from token skipping, the order of token logit calculation also matters a lot in computational cost. We apply a depth-first algorithm to traverse the token tree and calculate token logits so that the tokens that are calculated sequentially share the longest prefix sequence. This strategy reuses the KV cache of prefix sequences at most. Algorithm details are displayed in Appendix B.
@@ 4.3. Baselines
Table 2. Results of HtmlRAG without pruning and baselines under the long-context setting. Hit@1 is the proportion of instances where at least one short answer matches. The best and second best results are in bold and underlined. The symbol $\dagger$ signifies that our method achieves superior results among baselines in a statistically significant manner (t-test, $p$-value ¡ 0.05).
@@
Llama-3.1-8B-Instruct-128K
Vanilla HTML 47.75 20.08 28.75 47.25 36.09 85.00 24.85 6.00 16.13 6.28
Plain Text 61.50 27.82 39.25 59.25 44.31 94.00 28.23 7.75 16.02 6.35
Markdown 61.75 26.70 37.50 57.50 42.85 91.50 26.67 7.50 16.12 5.91
Llama-3.1-70B-Instruct-128K
@@ 4.4. Experimantal Settings
For a fair comparison, all end-to-end QA results are experimented with the latest open-source LLM, Llama-3.1-70B-Instruct and Llama-3.1-8B-Instruct (Dubey et al., 2024) under a 4K context window. As for the implementation details of our method, we construct a block tree with a granularity of 256 words before pruning with the embedding model, and we construct a finer-grained block tree with a granularity of 128 words before pruning with the generative model. We choose BGE-Large-EN (Xiao et al., 2023) as the embedding model for the HTML pruning. We choose a lightweight Phi-3.5-Mini-Instruct (Abdin et al., 2024) with 3B parameters as the backbone for our generative model. The training data used in fine-tuning the generative model contains 2635 automatically constructed training samples ranging from 2K to 32K in length. More implementation details can be found in Appendix A.
@@ 4.6.1. The Effectiveness of HTML Cleaning
The cleaned HTML is still long, so we conduct experiments under a long-context setting (128K), as shown in Table 2. When HTML is taken as the format of external knowledge, HtmlRAG without pruning meets or outperforms plain text and Markdown on most datasets, demonstrating its validity. Besides, we make the following observations: (1) Unprocessed HTML documents contain a large amount of irrelevant content, so all cleaning algorithms show improvements over vanilla HTML. (2) A more capable LLM (70B) performs better than a less capable one (8B) when taking HTML as the format of external knowledge. This indicates that more powerful models are better at understanding the complex information within HTML.
@@ 4.6.2. Ablation Study
Result Length # Params Storage # In-Tokens # Out-Tokens
BGE 200M 2.5G 93.54K 740.3
Prune-Embed 200M 2.5G 152.5K 2653
Prune-Gen 3B 7.2G 6750 28.70
LLM Chat 70B 131G 3661 182.9
@@ 5. Conclusion and Future Works
Appendix A Generative Model Training Details

### v2

@@ Abstract.
Retrieval-Augmented Generation (RAG) has been shown to improve knowledge capabilities and alleviate the hallucination problem of LLMs. The Web is a major source of external knowledge used in RAG systems, and many commercial RAG systems have used Web search engines as their major retrieval systems. Typically, such RAG systems retrieve search results, download HTML sources of the results, and then extract plain texts from the HTML sources. Plain text documents or chunks are fed into the LLMs to augment the generation. However, much of the structural and semantic information inherent in HTML, such as headings and table structures, is lost during this plain-text-based RAG process. To alleviate this problem, we propose HtmlRAG, which uses HTML instead of plain text as the format of retrieved knowledge in RAG. We believe HTML is better than plain text in modeling knowledge in external documents, and most LLMs possess robust capacities to understand HTML. However, utilizing HTML presents new challenges. HTML contains additional content such as tags, JavaScript, and CSS specifications, which bring extra input tokens and noise to the RAG system. To address this issue, we propose HTML cleaning, compression, and a two-step block-tree-based pruning strategy, to shorten the HTML while minimizing the loss of information. Experiments on six QA datasets confirm the superiority of using HTML in RAG systems. Our code and datasets are available at https://github.com/plageon/HtmlRAG.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the ACM Web Conference 2025; April 28–May 2, 2025; Sydney, NSW, Australia.booktitle: Proceedings of the ACM Web Conference 2025 (WWW ’25), April 28–May 2, 2025, Sydney, NSW, Australiaisbn: 979-8-4007-1274-6/25/04doi: 10.1145/3696410.3714546ccs: Information systems Web search engines
@@ 1. Introduction
However, employing HTML as the knowledge format for LLMs also presents the challenge of handling longer input sequences and noisy contexts. Our preliminary experiments show that a real HTML document from the Web contains over 80K tokens on average, among which over 90% of the tokens are CSS styles, JavaScript, Comments, or other meaningless tokens. Compared to the common maximum context window of current LLMs, which ranges from 32K to 128K, an individual document length of 80K is unacceptable. The aforementioned meaningless tokens in HTML documents can also affect the generation quality of LLMs. To solve this problem, in this paper, we devise a HTML Cleaning module to remove semantically irrelevant content in HTML documents, while keeping the main content intact. We also adjust the HTML tree structure without losing semantic information, for example, merging multiple layers of single nested HTML tags and removing empty tags. These processes reduce the length of the HTML to 6% of its original size.
Even after cleaning, HTML documents remain relatively long (over 4K each) to LLMs. To shorten the input context and remove the noise contained in the original retrieved documents, existing RAG systems have utilized different types of post-retrieval result refiners (Zhou et al., 2024b; Jin et al., 2024c; Xu et al., 2024; Jiang et al., 2024). These refiners extract the relevant text chunks or key sentences from the documents, regarding the user’s query and LLMs’ preference, and discard other content. These plain-text-based refiners cannot be directly applied to HTML because simply chunking HTML without considering its structure may generate unreasonable chunks. Hence, we further design an HTML Pruning module, which functions upon the intrinsic tree structure of HTML. The pruning process is comprised of the following steps:
@@ 2.1. Retrieval-Augmented Generation (RAG)
RAG systems augment LLM with external knowledge. A typical RAG pipeline includes components such as a query rewriter (Tan et al., 2024), a retriever (Li et al., 2024a; Shi et al., 2024; Cheng et al., 2024a; Mo et al., 2024), a reranker (Wang et al., 2024c; Shi et al., 2024), a refiner (Xu et al., 2024; Jin et al., 2024c; Jiang et al., 2024), and a reader (Zhu et al., 2024; Asai et al., 2024). This typical pipeline is widely used by mainstream RAG frameworks, such as LangChain (Chase, 2022) and LlamaIndex (Liu, 2022). Many works aim to optimize components in the pipeline, and previous works also manage to enhance the performance of RAG in other ways. Some methods devise new RAG frameworks, like retrieving external knowledge actively when internal knowledge is missing (Tan et al., 2024; Jiang et al., 2023b; Asai et al., 2024), or letting the LLM plan the retrieval process in a straight line or a tree structure (Shao et al., 2023; Kim et al., 2023b). However, most existing RAG systems take plain text as the format of external knowledge (Jin et al., 2024b; Dong et al., 2024; Cheng et al., 2024b; Li et al., 2024b). Instead, we propose to take HTML as a new format, and we believe using HTML can keep richer semantics in retrieved results.
@@ 2.2. Post-Retrieval Process of RAG
RAG systems usually apply post-retrieval processes (i.e., result refiners) to extract only the useful content to shorten the input context sent to LLMs. The chunking-based refiner is a widely used solution, which first chunks the text according to certain rules, and then uses a reranking model to select top chunks with high relevance (Karpukhin et al., 2020; Moniz et al., 2024; Li et al., 2025). Another solution is abstractive refiner, which utilizes a text-to-text language model to generate abstracts of results (Xu et al., 2024; Jiang et al., 2024; Gilbert et al., 2023). Some works use off-the-shelf abstractive models (Zhang et al., 2023a, b) or fine-tuned abstractive models (Jiang et al., 2024) to summarize retrieved results in a segmented and hierarchical manner. Others leverage the logits of language models to determine the importance of words within documents (Liu, 2019; Li, 2023).
@@
Figure 2. Overview of the HtmlRAG pipeline.
@@
Overview of the HtmlRAG pipeline.
@@ 3.2. HTML Cleaning
Since the original HTML documents are excessively long (over 80K each), and it’s needless to involve semantic features, model-based methods are inappropriate at this step. Thus, we first design a rule-based HTML cleaning, which pre-processes the HTML without considering the user’s query. This cleaning process removes irrelevant content and compresses redundant structures, retaining all semantic information in the original HTML. The compressed HTML afterof HTML cleaning is suitable for RAG systems equipped with long-context LLMs thatand are not willing to lose any information before generation. The cleaned HTML also serves as the basis for the following HTML pruning.
@@ 3.2.2. Lossless Structural Compression
3.3. Granularity-Adjustable Block Tree Building
@@
Considering the above problem, we propose an optimized tree structure that models HTML, which is not so fine-grained. Ideally, the granularity of the tree structure can be adjusted for different pruning requirements. We term it as a “block tree”, and we set the maximum number of words per block, $maxWords$ to control the granularity of the block tree. In terms of block tree construction, we start from a DOM tree, and we merge fragmented child nodes into their parent and treat them as a block. We can recursively merge blocks or child nodes into their parent to form a bigger block under the condition that the number of words in a block does not exceed $maxWords$. After merging, original leaf nodes that are unable to be merged are also regarded as blocks. Algorithm details are demonstrated in Appendix F.
@@ 3.4.1. Pruning Blocks based on Text Embedding
The refining process is expected to shorten the retrieval results while preserving key information as much as possible. A straightforward idea is to extract plain text in the block and calculate a similarity score with the user’s query using text embeddings. Then we use a greedy algorithm to prune the block tree by deleting low-similarity blocks and retraining higher ones. In practice, we keep deleting the block with the lowest relevance until the total length of the HTML documents satisfies the context window we set. After block deleting, redundant HTML structures will re-appear, so we re-adjust the HTML structure, meaning multiple layers of single-nested tags are merged and empty tags are removed. The detailed pruning algorithm is demonstrated in Appendix F.
@@ 3.4.2. Generative Fine-Grained Block Pruning
To further prune blocks with a finer granularity, we expand the leaf nodes of the pruned block tree and get a finer-grained block tree. Given the limitations of the embedding-model-based block pruning, we propose to use a generative model because it has a long context to cover the whole block tree and is not limited to modeling one block at a time. Yet processing the cleaned HTML directly with a generative model is inappropriate because the cleaned HTML is long (60K on average), which brings much computational cost. Similarly, the generative model is supposed to calculate scores for blocks. Inspired by CFIC (Qian et al., 2024) and Generative Retrieval (Li et al., 2024d, c), which takes the text chunk’s sequence generation probability as the score for that chunk, we propose to use a sequence of tags to identify a block. Specifically, the sequence consists of tags starting from the root tag and walking down to the block’s tag, and we term this sequence as “block path”. In the inference phase, the generative model follows the structure of the block tree and calculates the scores of blocks in the block tree. The scores of blocks are derived from the token logits, as displayed in Figure 3. At last, we use the same block pruning operation as we mention in §3.4.1 to obtain the refined HTML document.
@@
We collect a small amount of supervised data to enhance the model’s capability in block path generation. Following the typical SFT process (Qin et al., 2024), the steps for training data collecting, filtering, and constructing are as follows: First, we sample queries from the training set of several open-source QA datasets. For each query, we retrieve a couple of related HTML documents using the online search engine Bing. Then we clean the retrieved HTML, and prune the HTML with the embedding model. By adjusting the output length in HTML pruning, we get pruned HTML documents of various lengths, ranging from 2K tokens to 32K tokens. After that, we build a block tree from each HTML document pruned by the embedding model, and calculate the exact match score for the content within blocks with the gold answer. To ensure the data quality, we discard samples in which no block’s content exactly matches the gold answer, meaning highly relevant HTML documents are not retrieved. More training details are discussed in Appendix B.
@@
In the first two conditions, it is needless to infer with the generative model, meaning many tokens can be skipped. This brings down the inference computational cost. Apart from token skipping, the order of token logit calculation also matters a lot in computational cost. We apply a depth-first algorithm to traverse the token tree and calculate token logits so that the tokens that are calculated sequentially share the longest prefix sequence. This strategy reuses the KV cache of prefix sequences at most. Algorithm details are displayed in Appendix F.
@@ 4.3. Baselines
Table 2. Results of HtmlRAG without pruning and baselines under Llama-3.1-70B-Instruct-128K. Hit@1 is the proportion of instances where at least one short answer matches. The best and second best results are in bold and underlined. The symbol $\dagger$ signifies that our method achieves superior results among baselines in a statistically significant manner (t-test, $p$-value ¡ 0.05).
@@ 4.4. Experimantal Settings
For a fair comparison, all end-to-end QA results are experimented with the latest open-source LLM, Llama-3.1-70B-Instruct and Llama-3.1-8B-Instruct (Dubey et al., 2024) under a 4K context window. As for the implementation details of our method, we construct a block tree with a granularity of 256 words before pruning with the embedding model, and we construct a finer-grained block tree with a granularity of 128 words before pruning with the generative model. We choose BGE-Large-EN (Xiao et al., 2023) as the embedding model for the HTML pruning. We choose a lightweight Phi-3.5-Mini-Instruct (Abdin et al., 2024) with 3B parameters as the backbone for our generative model. The training data used in fine-tuning the generative model contains 2635 automatically constructed training samples ranging from 2K to 32K in length. More implementation details can be found in Appendix B.
@@ 4.6.1. The Effectiveness of HTML Cleaning
The cleaned HTML is still long, so we conduct experiments under a long-context setting (128K), as shown in Table 2. When HTML is taken as the format of external knowledge, HtmlRAG without pruning meets or outperforms plain text and Markdown on most datasets, demonstrating its validity. Besides, we make the following observations: (1) Unprocessed HTML documents contain a large amount of irrelevant content, so all cleaning algorithms show improvements over vanilla HTML. (2) Under a limited context window, HTML format reference contains less documents and has lower exact match score due to extra HTML tags occupying tokens. Under such circumstances, HTML performs comparable or even better than plain text. This shows the positive effect of the rich structural information of HTML.
@@ 4.6.2. Ablation Study
Method # Params Storage # In-Tokens # Out-Tokens Latency
BGE 200M 2.5G 93.54K 740.3 10.5ms
Prune-Embed 200M 2.5G 152.5K 2653 30.8ms
Prune-Gen 3B 7.2G 6750 28.70 6.15ms
LLM Chat 70B 131G 3661 182.9 915ms
@@ 5. Conclusion and Future Works
Acknowledgements.
Zhicheng Dou is the corresponding author. This work was supported by the National Natural Science Foundation of China No. 62272467, Beijing Natural Science Foundation No. L233008, Beijing Municipal Science and Technology Project No. Z231100010323009, the fund for building world-class universities (disciplines) of Renmin University of China. The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, MOE.
Appendix
Table 5. The information loss during pruning measured by the reference text’s exact match (EM) scores. The information loss of baselines are also demonstrated for comparison.
Length 128k 8k 4k
Dataset Plain Text Markdown HtmlRAG w/o Prune HtmlRAG w/o Gen BM25 BGE E5 - Mistral LongLLMLingua JinaAI Reader HtmlRAG
ASQA 70.90 67.72 69.49 59.72 25.99 52.17 50.65 35.77 38.39 52.80
HQA 70.00 65.75 66.75 54.75 37.50 48.75 49.00 43.75 38.25 51.00
Table 6. Complementary results of HtmlRAG without pruning and baselines under Llama-3.1-8B-Instruct-128K. Hit@1 is the proportion of instances where at least one short answer matches. The best and second best results are in bold and underlined. The symbol $\dagger$ signifies that our method achieves superior results among baselines in a statistically significant manner (t-test, $p$-value ¡ 0.05).
Method ASQA Hotpot-QA NQ Trivia-QA MuSiQue ELI5
Hit@1 EM EM Hit@1 EM Hit@1 EM EM ROUGE-L BLEU
Vanilla HTML 47.75 20.08 28.75 47.25 36.09 85.00 24.85 6.00 16.13 6.28
Plain Text 61.50 27.82 39.25 59.25 44.31 94.00 28.23 7.75 16.02 6.35
Markdown 61.75 26.70 37.50 57.50 42.85 91.50 26.67 7.50 16.12 5.91
Appendix A Generalization to Rich Text Formats
The HTML cleaning and pruning process is worthwhile as long as the knowledge source is in HTML format or in other rich formats like PDF. Currently, major RAG frameworks like LangChain and LlamaIndex share the following workflow: “Retrieve HTML -¿ Convert to Plain Text -¿ Refine -¿ Generate Answer”. We argue that the upper bound of the workflow above is limited because lots of information is lost during the early HTML to plain text conversion. Our proposed workflow goes like this: “Retrieve HTML -¿ Clean and Prune -¿ Convert to Other Formats (Optional) -¿ Generate Answer”. Even if the LLM prefers other input formats or you’d like to save tokens, the conversion from HTML to other formats is optional and suggested to be done after HTML cleaning and pruning. This workflow has a higher upper bound because pruning is carried out before the information loss of format conversion.
Figure 6. Node content explained
[…]
HTML for RAG pipeline overview
Appendix B Generative Model Training Details